OpenStudy (anonymous):

I need to find the limit as x goes to 9. I must use factoring. How can I factor? ((x^(3/2))-27) / x-9

myininaya (myininaya):

\[(x^\frac{1}{2})^3-3^3=(x^\frac{1}{2}-3)(x+3x^\frac{1}{2}+9)\] You can look at the stop as a difference of cubes You could look at the bottom as a difference of squares and factor and see something cancels